OpenStudy (anonymous):

Hi Given the following system of linear equations; ax + y + z = 1, x + ay + z = a, x + y + az = a^2 find the values of the real constant, a, for which the system has a unique solution. Thanks for the help.

OpenStudy (unklerhaukus):

a=1

OpenStudy (blockcolder):

Lemme try this: Add all three equations: (a+2)x+(a+2)y+(a+2)z=a^2+a+1 (x+y+z)(a+2)=a^2+a+1 x+y+z=(a^2+a+1)/(a+2) Now subtract this from your equations.

OpenStudy (unklerhaukus):

has 3 fold infinite solution

OpenStudy (anonymous):

Ahh thanks.

OpenStudy (unklerhaukus):

a= 0 has a single solution